Suppose that at least 17.8 eV is needed to free an electron from a particular element, i.e., to ionize the atom. What is the lowest frequency photon which can accomplish this? And, what is the corresponding wavelength for that photon? Give frequency for answer 1 and wavelength for answer 2.
Answer #1

energy of photon, frequency (17.8eV)(1.6x10-19 T R leV v=-= 6.63x10-43x101Hz wavelength c = 3,0 x 108 m/s = 700x 10-1m = 700